# Robust Policy Switching for Antifragile Reinforcement Learning for UAV Deconfliction in Adversarial Environments

### v1

@@ IV Discounted Thompson Sampling for Switching Robust Policies
Figure 4: Discounted Thompson sampling (DTS) technique to select the for antifragility.
@@
As illustrated in Figure 4, the antifragility framework employs a DTS-based mechanism to dynamically switch between robust policies in response to adversarial observations. Initially, adversarial observations, denoted as $\Phi_{\textup{adv}}$, are synthesized from the original observations $\Phi$ by incorporating the gradient sign of DDPG actor loss, $\nabla_{\Phi}\mathcal{L}_{\pi}$, with the maximum magnitude of the perturbation $\epsilon$. Using the robust action policy, as described in the previous section, the corresponding adversarial action $\mathbf{a}_{\textup{adv}}$ is derived. Currently, actions $\mathbf{a}$ are obtained from the original observations $\Phi$ using the deep deterministic policy gradient (DDPG) policy. To assess the impact of adversarial observations, the value distributions are obtained from trained critics, specifically $Q(\Phi,\mathbf{a})$ for vanilla DDPG and $Q_{\textup{rob}}(\Phi_{\textup{adv}},\mathbf{a}_{\textup{adv}})$ for action robust DDPG. The distributional shift is quantified using the Wasserstein distance between these value functions. Subsequently, the distributional shift is normalized with respect to the minimum distance in all robust value functions to compute the true probability ${p}_{\textup{true}}$.
@@
Given the robust model selected, the Bernoulli reward, which represents the likelihood, is calculated based on a Beta prior with parameters $(\mathfrak{a},\mathfrak{b})$, while updating the conjugate Beta posterior distribution. After a certain number of time steps, once the posterior update for the DTS is complete, the attack strength is increased by $\Delta\epsilon$, resulting in a new set of probabilities, $\mathbf{p}_{\textup{true}}$, thus addressing the nonstationary Bernoulli reward. Consequently, the DTS parameters $(\mathfrak{a},\mathfrak{b})$ are updated with a discount factor $\mathfrak{y}$, which accounts for the temporal variations in $\mathbf{p}_{\textup{true}}$. The methodology for generating adversarial observations using the Frank-Wolfe optimization technique is presented in the following subsection.
IV-A Adversarial Observations
@@
In the context of antifragility, adaptation must take into account adversarial observations to ensure robustness. Given that the structure of the model is known, adversarial perturbations are formulated under a white-box attack framework. In this paper, a projection-free adversarial attack mechanism based on the Frank-Wolfe algorithm [chen2020frank] is used. This approach leverages an iterative first-order white-box attack with momentum, which exhibits a convergence rate of $O\left(1\mathbin{/}\sqrt{T}\right)$, where $T$ represents the number of steps. For a given state representation $\Phi$, let $\mathcal{P}_{\mathcal{X}}\left({\Phi}\right)$ denote the projection operation that maps the observation $\Phi$ to the constraint set $\mathcal{X}$ at each iteration. In contrast to the projected gradient descent attack (PGD) [gupta2018cnn], the Frank-Wolfe algorithm is inherently projection-free, as it instead relies on a linear minimization oracle (LMO) over the constraint set $\mathcal{X}$ at each iteration. The LMO is defined as follows,
@@
The minimization procedure to obtain an adversarial state in (13) can be implemented using the process shown in Algorithm 2. In the Frank-Wolfe algorithm, the linear minimization oracle (LMO) optimally solves a linear subproblem in the constraint set,$\mathcal{X}$, followed by a weighted averaging step with the previous iterate to compute the final update as shown in Algorithm 2. Compared to the PGD attack, the Frank-Wolfe approach is more conservative, as it eliminates the need for explicit projection steps, resulting in improved distortion characteristics.
@@ IV-B Distributional Shift Due to Adversarial Observation
The shift in the value distribution caused by adversarial observations is quantified using the 1-Wasserstein norm, which is normally used to measure the distance between probability distributions. Let $\mathcal{P}\left(\mathbb{R}^{d}\right)$, denote the space of probability measures in $\mathbb{R}^{d}$ with finite first moments. Given two probability measures, $\beth,\daleth\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, the 1-Wasserstein is defined as:
@@
Here $\Pi\left(\beth,\daleth\right)$ represents the set of all joint distributions $\iota$ with marginals $\beth$ and $\daleth$. $\left\|x-y\right\|$ represents the Euclidean norm in $\mathbb{R}^{d}$, and $\mathbb{E}_{\left(x,y\sim\iota\right)}\left[\left\|x-y\right\|\right]$ quantifies the cost of transporting mass from $x$ to $y$. Similarly, if we consider a critic for vanilla DDPG be $Q$ and for action robust critic as $Q^{\alpha}_{\textup{AR}}$, for which we can define the value distribution $Z\left(\Phi\right)$ and $Z^{\alpha}_{\textup{AR}}\left(\Phi_{\textup{adv}}\right)$ follows:
$\begin{gathered}Z\left(\Phi\right)=Q_{\textup{DDPG}}\left(\Phi,\pi\left(\Phi\right)\right),\\ Z^{\alpha}_{\textup{AR}}\left(\Phi\right)=Q^{\alpha}_{\textup{AR}}\left(\Phi_{\textup{adv}},\alpha\cdot\pi_{\textup{adv}}\left(\Phi\right)+\left(1-\alpha\right)\cdot\pi_{\textup{agent}}\right).\\ \end{gathered}$ (15)
We define the distribution shift $d^{\alpha}_{\epsilon}$ due to adversarial observation $\Phi_{\textup{adv}}$ due to adversarial attack of intensity $\epsilon$ as,
@@
Similarly for DDPG policy,$\pi_{\textup{DDPG}}$, we define the value distribution at adversarial observation $\Phi_{\textup{adv}}$ as,
@@
According to (20), a lower value of $d^{\alpha}_{\epsilon}$ for a given robust model indicates a smaller shift in the state-action value distribution under an adversarial attack of magnitude $\epsilon$. Consequently, a higher value $p_{\textup{true}}$ for a specific model enhances its likelihood of successful deployment in adversarial settings.
@@ IV-C Discounted Thompson Sampling for Minimum Distribution Shift
Let $\mathcal{K}=\left\{\pi_{0},\pi_{1},\cdots,\pi_{n}\right\}$ represent the set of robust policy networks as multi-armed bandits (MAB) available to counter an adversarial observation. Given a finite time horizon $T$, the DTS must select a policy network $\pi_{k}\in\mathcal{K}$ at each time step $t\in\left\{1,\cdots,T\right\}$. Let $X_{k,t}$ denote the reward obtained by employing the $k$-th robust policy at time $t$. Consequently, the instantaneous reward $X_{k,t}$ follows a Bernoulli distribution with an expected value of $\grave{\mu}_{k,t}=\mathbb{E}\left[X_{k,t}\right]$. Therefore, the optimal policy at any given time $t$ is the one that maximizes the expected reward, expressed as $\mathbf{\grave{\mu}}^{*}_{t}=\max_{k\in\mathcal{K}}\left\{\grave{\mu}_{k,t}\right\}$.
In non-stationary environments, the expected Bernoulli rewards evolve over time, either following a gradual trend or undergoing abrupt changes. The sequence of rewards for a specific policy $\pi_{k}$ is denoted by $\grave{\mu}^{k}=\left\{\grave{\mu}_{k,t}\right\}_{t=1}^{T}$, while the general sequence of rewards across all policies is represented as $\grave{\mu}=\left\{\grave{\mu}^{k}\right\}_{k\in\mathcal{K}}$. If we consider $\mathcal{P}$ to be the family of admissible antifragile policies for DTS, let $\varkappa\in\mathcal{P}$ represent a candidate policy responsible for selecting an appropriate robust policy in response to adversarial observations. At each time step $t$, the policy $\varkappa$ chooses the robust policy $I^{\varkappa}_{t}$ based on an initial prior $U$ and past switching decisions $\left\{X_{I^{\varkappa}_{t},n}\right\}_{n=1}^{t-1}$. Hence, the DTS switching policy $\varkappa_{t}$ at time $t$ is given by,
$\varkappa_{t}=\begin{cases}\varkappa_{1}(U)&;\,t=1\\ \varkappa_{t}(U,X_{I_{\varkappa_{1},1}},\ldots,X_{I_{\varkappa_{t-1},t-1}})&;\,t\geq 2\end{cases}.$ (21)
Thompson Sampling (TS) [agrawal2012analysis] operates by maintaining a prior distribution over the success probabilities of each Bernoulli-based robust policy and sampling from this prior to determine which robust model to deploy. The prior distribution is typically modeled as a Beta distribution $\mathfrak{B}\left(\mathfrak{a},\mathfrak{b}\right)$, with the parameters $\left\{\mathfrak{a},\mathfrak{b}\right\}$ updated based on the Bernoulli rewards obtained from the deployment of the robust model against adversarial observation.
@@
$\begin{gathered}\mathbb{E}\mathfrak{S}_{k,t+1}=\mathfrak{y}\mathbb{E}\mathfrak{S}_{k,t}+\breve{\mu}_{k,t}\mathbb{P}\left(I^{\varkappa}_{t}=k\right),\\ \mathbb{E}\mathfrak{F}_{k,t+1}=\mathfrak{y}\mathbb{E}\mathfrak{F}_{k,t}+\left(1-\breve{\mu}_{k,t}\right)\mathbb{P}\left(I^{\varkappa}_{t}=k\right).\end{gathered}$ (22)
@@
$\breve{\sigma}^{2}_{k,t+1}=\frac{\mathfrak{S}_{k,t+1}\cdot\mathfrak{F}_{k,t+1}}{(\mathfrak{S}_{k,t+1}+\mathfrak{F}_{k,t+1})^{2}(\mathfrak{S}_{k,t+1}+\mathfrak{F}_{k,t+1}+1)},$ (24)
$=\frac{\breve{\mu}_{k,t}(1-\breve{\mu}_{k,t})}{\mathfrak{y}\cdot\mathfrak{S}_{k,t}+\mathfrak{y}\cdot\mathfrak{F}_{k,t}+1}\geq\frac{\breve{\mu_{k,t}}(1-\breve{\mu_{k,t}})}{\mathfrak{S}_{k,t}+\mathfrak{F}_{k,t}+1}.$ (25)
@@ V Theoretical Analysis
In this section we will theoretically prove that DTS based switching of robust policies, help the algorithm deal with adaptation against unseen attacks and induce antifragility.
@@ V-A Regret Minimization
Above assumption takes into account the engineering realism of non-sudden shift in the reward due to adversarial states. Let us define a proxy reward to assess robustness, similar to (20), which assigns higher reward to policies which exhibits smaller value shifts under perturbation.
@@
It shows that minimizing the regret is equivalent to selecting the DTS policy with the minimal value shift which justifies the use of $d^{\left(t\right)}_{k}$ in (19) as a robustness metric.
As shown in [wu2018learning], DTS achieves sublinear regret in multi-armed bandit problems with $M$ changepoints, shown as
@@
Substituting into our bound, we get,
@@
where $\kappa_{t}$ is the policy selected by DTS at time t. As the regret is minimized with respect to non-stationary adversarial attacks, now we see the nature of the algorithm against unseen adversarial attacks.
V-B Adaption to Unseen Adversarial Attacks
@@
$\lim_{t\to\infty}\mathbb{P}(\kappa_{t}=k^{*})=1.$
@@
$\tilde{r}_{k}^{(t)}\sim\text{Beta}(S_{k}^{(t)}+a_{0},F_{k}^{(t)}+b_{0}).$
@@
$S_{k}^{(t)}=\sum_{i=1}^{t}\gamma^{t-i}r_{k}^{(i)},\quad F_{k}^{(t)}=\sum_{i=1}^{t}\gamma^{t-i}(1-r_{k}^{(i)}).$
@@
$\mathbb{P}(\tilde{r}_{j}^{(t)}>\tilde{r}_{k^{*}}^{(t)}).$
@@
$\mathbb{P}(\tilde{r}_{j}^{(t)}>\tilde{r}_{k^{*}}^{(t)})\leq\exp\left(-\frac{(\mu_{k^{*}}-\mu_{j})^{2}}{C(1-\gamma)^{2}}\right),$
@@
$\boxed{\lim_{t\to\infty}\mathbb{P}(\kappa_{t}=\arg\min_{k}d_{k}^{(t)})=1}$
@@
Theorem 1 implies that the DTS mechanism enables robust policy adaptation even in the presence of previously unseen or out-of-distribution (OOD) adversarial perturbations. Although the policy ensemble $\mathcal{P}=\{\pi_{\alpha_{1}},\dots,\pi_{\alpha_{K}}\}$ is trained only on a discrete set of known adversarial intensities $\{\alpha_{k}\}$, DTS does not require explicit knowledge of the true test-time perturbation $\epsilon_{t}\notin\{\alpha_{1},\dots,\alpha_{K}\}$. Instead, DTS relies on online estimates of the value distribution shift $d_{k}^{(t)}$ as a proxy for robustness. As long as one policy in $\mathcal{P}$ exhibits relatively low Wasserstein shift under the unseen attack (i.e., approximates the behavior needed to withstand $\epsilon_{t}$), its reward proxy $r_{k}^{(t)}=d_{\min}^{(t)}/d_{k}^{(t)}$ will dominate over time. Due to posterior concentration, DTS will asymptotically prefer this policy, allowing the agent to adapt dynamically to new adversarial conditions without retraining. This property ensures that the agent is not merely robust, but adaptive to distributional shifts caused by novel adversarial strategies.

### v3

@@ IV Discounted Thompson Sampling for Switching Robust Policies
Figure 3: Meta-policy design using Discounted Thompson sampling (DTS) to select the robust policies against self-induced adversarial observations.
As illustrated in Figure 3, we present the novel DTS-based adaptive policy-switching framework for adaptive robustness against unseen adversarial attacks. This method uniquely addresses the limitation of the existing adaptive adversarial RL algorithms [pinto2017robust, rakelly2019efficient] ,which are computationally expensive and impractical in real-time UAV operations, as it dynamically selects among pre-trained robust policies using discounted DTS based on real-time estimation of the value distribution shifts thus offering efficient, real-time adaptive robustness against adversarial perturbations without any retraining during deployment.
@@
Adversarial observations, denoted as $\Phi_{\textup{adv}}$, are synthesized from the original observations $\Phi$ by incorporating the gradient sign of DDPG actor loss, $\nabla_{\Phi}\mathcal{L}_{\pi}$, with the maximum magnitude of the perturbation $\epsilon$. Using the robust action policy, as described in the previous section, the corresponding adversarial action $\mathbf{a}_{\textup{adv}}$ is obtained. Similarly, from the nominal policy $\pi$ and original observations $\Phi$, the agent actions $\mathbf{a}$ are obtained. To assess the impact of adversarial observations, the value distributions are obtained from trained critics, specifically $Q(\Phi,\mathbf{a})$ for vanilla DDPG and $Q_{\textup{rob}}(\Phi_{\textup{adv}},\mathbf{a}_{\textup{adv}})$ for action robust DDPG. The distributional shift is quantified using the Wasserstein distance between these value functions. Subsequently, the distributional shift is normalized with respect to the minimum distance for all robust value functions to compute the true probability ${p}_{\textup{true}}$.
@@
With the selected robust policy, the Bernoulli reward, which represents the likelihood, is calculated based on a Beta prior with parameters $(\mathfrak{a},\mathfrak{b})$, while updating the conjugate Beta posterior distribution. After a certain number of time steps, once the posterior update for the DTS is complete, the self-induced attack strength is increased by $\Delta\epsilon$, resulting in a new set of probabilities, $\mathbf{p}_{\textup{true}}$, thus addressing the nonstationary Bernoulli reward. Consequently, the DTS parameters $(\mathfrak{a},\mathfrak{b})$ are updated with a discount factor $\mathfrak{y}$, which accounts for the temporal variations in $\mathbf{p}_{\textup{true}}$. The methodology for generating self-induced adversarial observations using the Frank-Wolfe optimization technique is presented in the following subsection.
IV-A Self-Induced Adversarial Observations
@@
Here, the meta-policy switching adaptation is obtained using self-induced adversarial observations to ensure robustness. The adversarial observations are called self-induced in this context, as it is generated from nominal policy $\pi$ known to the agent. Here, a projection-free self-induced adversarial attack mechanism based on the Frank-Wolfe algorithm [chen2020frank] is used. This approach leverages an iterative first-order white-box attack with momentum, which exhibits a convergence rate of $O\left(1\mathbin{/}\sqrt{T}\right)$, where $T$ represents the number of steps. For a given state representation $\Phi$, let $\mathcal{P}_{\mathcal{X}}\left({\Phi}\right)$ denote the projection operation that maps the observation $\Phi$ to the constraint set $\mathcal{X}$ at each iteration. In contrast to the projected gradient descent attack (PGD) [gupta2018cnn], the Frank-Wolfe algorithm is inherently projection-free, as it instead relies on a linear minimization oracle (LMO) over the constraint set $\mathcal{X}$ at each iteration. The LMO is defined as follows,
@@
[…]
@@ IV-B Distributional Shift Due to Adversarial Observation
The shift in the value distribution caused by self-induced adversarial observations is quantified using the 1-Wasserstein norm, which is normally used to measure the distance between probability distributions. Let $\mathcal{P}\left(\mathbb{R}^{d}\right)$ denote the space of probability measures in $\mathbb{R}^{d}$ with finite first moments. Given two probability measures, $\beth,\daleth\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, the 1-Wasserstein is defined as:
@@
[…]
We define the distribution shift $d^{\alpha}_{\epsilon}$ due to adversarial observation $\Phi_{\textup{adv}}$ with the intensity $\epsilon$ as,
@@
[…]
@@
According to (20), a lower value of $d^{\alpha}_{\epsilon}$ for a given robust model indicates a smaller shift in the state-action value distribution under a self-induced adversarial observation. Consequently, a higher value $p_{\textup{true}}$ for a specific model enhances its likelihood of successful deployment in adversarial settings. Given that $d_{\text{min}}<d^{\alpha}_{\epsilon}$ , we can expect that $\mathbf{p}_{\text{true}}$ is bounded within $\left(0,1\right)$ if we consider $k_{\text{mult}}\in\left(0,1\right)$. The inverse relationship between $d^{\alpha}_{\epsilon}$ and $\mathbf{p}_{\text{true}}$ enables the DTS to prefer policies that induce minimal value distribution shift under attack, thus enhancing adaptive resilience.
@@ IV-C Discounted Thompson Sampling for Minimum Distribution Shift
Let $\mathcal{K}=\left\{\pi_{0},\pi_{1},\cdots,\pi_{n}\right\}$ represent the ensemble set of robust policies as multi-armed bandits (MAB) available to counter an adversarial observation. Given a finite time horizon $T$, the DTS must select a policy network $\pi_{k}\in\mathcal{K}$ at each time step $t\in\left\{1,\cdots,T\right\}$. Let $X_{k,t}$ denote the reward obtained by employing the $k$-th robust policy at time $t$. Consequently, the instantaneous reward $X_{k,t}$ follows a Bernoulli distribution with an expected value of $\grave{\mu}_{k,t}=\mathbb{E}\left[X_{k,t}\right]$. Therefore, the optimal policy at any given time $t$ is the one that maximizes the expected reward, expressed as $\mathbf{\grave{\mu}}^{*}_{t}=\max_{k\in\mathcal{K}}\left\{\grave{\mu}_{k,t}\right\}$.
In non-stationary environments, the expected Bernoulli rewards evolve over time, either following a gradual trend or undergoing abrupt changes. The sequence of rewards for a specific policy $\pi_{k}$ is denoted by $\grave{\mu}^{k}=\left\{\grave{\mu}_{k,t}\right\}_{t=1}^{T}$, while the general sequence of rewards across all policies is represented as $\grave{\mu}=\left\{\grave{\mu}^{k}\right\}_{k\in\mathcal{K}}$. If we consider $\mathcal{P}$ to be the family of admissible policies for DTS, let $\varkappa\in\mathcal{P}$ represent a candidate policy responsible for selecting an appropriate robust policy in response to adversarial observations. At each time step $t$, the policy $\varkappa$ chooses the robust policy $I^{\varkappa}_{t}$ based on an initial prior $U$ and past switching decisions $\left\{X_{I^{\varkappa}_{t},n}\right\}_{n=1}^{t-1}$. Hence, the DTS switching policy $\varkappa_{t}$ at time $t$ is given by,
$\varkappa_{t}=\begin{cases}\varkappa_{1}(U)&;\,t=1\\ \varkappa_{t}(U,X_{I_{\varkappa_{1},1}},\ldots,X_{I_{\varkappa_{t-1},t-1}})&;\,t\geq 2.\end{cases}$ (21)
Thompson Sampling (TS) [agrawal2012analysis] operates by maintaining a prior distribution over the success probabilities of each Bernoulli-based robust policy and sampling from this prior to determine which robust model to deploy. The prior distribution is typically modeled as a Beta distribution $\mathfrak{B}\left(\mathfrak{a},\mathfrak{b}\right)$, with the parameters $\left\{\mathfrak{a},\mathfrak{b}\right\}$ updated based on the Bernoulli rewards obtained from the deployment of the robust model against adversarial observations.
@@
$\begin{gathered}\mathbb{E}\left[\mathfrak{S}_{k,t+1}\right]=\mathfrak{y}\mathbb{E}\left[\mathfrak{S}_{k,t}\right]+\breve{\mu}_{k,t}\mathbb{P}\left(I^{\varkappa}_{t}=k\right),\\ \mathbb{E}\left[\mathfrak{F}_{k,t+1}\right]=\mathfrak{y}\mathbb{E}\left[\mathfrak{F}_{k,t}\right]+\left(1-\breve{\mu}_{k,t}\right)\mathbb{P}\left(I^{\varkappa}_{t}=k\right).\end{gathered}$ (22)
@@
$\begin{gathered}\breve{\sigma}^{2}_{k,t+1}=\frac{\mathfrak{S}_{k,t+1}\cdot\mathfrak{F}_{k,t+1}}{(\mathfrak{S}_{k,t+1}+\mathfrak{F}_{k,t+1})^{2}(\mathfrak{S}_{k,t+1}+\mathfrak{F}_{k,t+1}+1)},\\ =\frac{\breve{\mu}_{k,t}(1-\breve{\mu}_{k,t})}{\mathfrak{y}\cdot\mathfrak{S}_{k,t}+\mathfrak{y}\cdot\mathfrak{F}_{k,t}+1}\geq\frac{\breve{\mu_{k,t}}(1-\breve{\mu_{k,t}})}{\mathfrak{S}_{k,t}+\mathfrak{F}_{k,t}+1}.\end{gathered}$ (24)
@@
IV-D Test-Time Adversarial Attacks
To evaluate the adaptive robustness of the DTS-based policy switching mechanism, we test it against two distinct and practically motivated classes of adversarial threats: (1) observation-space adversarial attacks induced via gradient-based perturbations, and (2) signal-level attacks via GNSS spoofing, which introduce physical misalignments in the UAV’s perceived position. This dual-threat framework captures both control-layer and sensor-layer vulnerabilities relevant to dependable cyber-physical systems. The practicality of test-time threats is elaborated in Appendix B.
@@ V Theoretical Analysis
In this section we will theoretically prove that DTS based robust policy switching can ensure adaptive robustness by minimizing the regret and also ensure antifragile behaviour.
@@ V-A Regret Minimization
The assumption stated in (25) takes into account the engineering realism of non-sudden shift in the reward due to self-induced adversarial states. Let us define a proxy reward to assess robustness, similar to (20), which assigns higher reward to policies which exhibits smaller value shifts under perturbation.
@@
The relation in (28) shows that minimizing the regret is equivalent to selecting the DTS policy with the minimal value shift which justifies the use of $d^{\left(t\right)}_{k}$ in (19) as a robustness metric. As shown in [wu2018learning], DTS achieves sublinear regret in multi-armed bandit problems with $M$ changepoints, shown as
@@
This analysis is analogous as the changepoints represent the time-instants where the adversarial attack strength was increased. Substituting into our bound in (29), we get,
@@
where $\kappa_{t}$ is the policy selected by DTS at time $t$. As the regret is to be minimized with respect to non-stationary self.induced adversarial observations, now we highlight the adaptive robustness and antifragile behaviour of the meta-policy switching framework.
V-B Adaptive Robustness against Unseen Adversarial Attacks
@@
$\lim_{t\to\infty}\mathbb{P}(\kappa_{t}=k^{*})=1.$ (31)
@@
$\tilde{r}_{k}^{(t)}\sim\text{Beta}(S_{k}^{(t)}+a_{0},F_{k}^{(t)}+b_{0}).$ (32)
@@
$S_{k}^{(t)}=\sum_{i=1}^{t}\gamma^{t-i}r_{k}^{(i)},\quad F_{k}^{(t)}=\sum_{i=1}^{t}\gamma^{t-i}(1-r_{k}^{(i)}).$ (33)
@@
$\mathbb{P}(\tilde{r}_{j}^{(t)}>\tilde{r}_{k^{*}}^{(t)}).$ (34)
@@
$\mathbb{P}(\tilde{r}_{j}^{(t)}>\tilde{r}_{k^{*}}^{(t)})\leq\exp\left(-\frac{(\mu_{k^{*}}-\mu_{j})^{2}}{C(1-\gamma)^{2}}\right),$ (35)
@@
$\boxed{\lim_{t\to\infty}\mathbb{P}(\kappa_{t}=\arg\min_{k}d_{k}^{(t)})=1}$ (37)
@@
Theorem 1 implies that the DTS mechanism enables robust policy adaptation even in the presence of previously unseen or out-of-distribution (OOD) adversarial perturbations. Although the policy ensemble $\mathcal{P}=\{\pi_{\alpha_{1}},\dots,\pi_{\alpha_{K}}\}$ is trained only on a discrete set of known self-induced adversarial intensities $\{\alpha_{k}\}$, DTS does not require explicit knowledge of the true test-time perturbation $\epsilon_{t}\notin\{\alpha_{1},\dots,\alpha_{K}\}$. Instead, DTS relies on online estimates of the value distribution shift $d_{k}^{(t)}$ as a proxy for robustness. As long as one policy in $\mathcal{P}$ exhibits relatively low Wasserstein shift under the unseen attack (i.e., approximates the behavior needed to withstand $\epsilon_{t}$), its reward proxy $r_{k}^{(t)}=d_{\min}^{(t)}/d_{k}^{(t)}$ will dominate over time. Due to posterior concentration, DTS will asymptotically prefer this policy, allowing the agent to adapt dynamically to new adversarial conditions without retraining. This property ensures that the agent is not merely robust, but adaptive to distributional shifts caused by novel adversarial strategies.
Remark 3.
The theoretical analysis in this work is grounded on several key assumptions, notably the Lipschitz continuity of the reward function under the value distribution shifts, and slowly-varying nature of self-induced adversarial perturbations to obtain the DTS policy. These assumptions, which are standard in the literature for achieving robustness [song2023lipsnet] , support in obtaining piecewise stationary DTS policy and also analyzing regret during test-time attacks. These assumptions are valid as they are based on slow reward change due adversarial observations which are self-induced. The assumptions are used merely to learn the adaptation, so it can deal with persistent spoofing threats which are OOD in nature.